21-Aug-2017 02:00 GMT to 21-Aug-2017 06:00 GMT
During the above window one of our fibre suppliers will be carrying out maintenance on one of our links between Kingston and Slough. Whilst the actual downtime is expected to be less, the full 4 hrs may be required to account for any unforeseen issues. This work is not expected to be service affecting, however, you may see unusual paths and increase in latency while traffic is automatically rerouted via alternate routes.
We are currently experiencing a loss of connectivity to one of our switches in Interxion. Engineers are already investigating and local engineering resources are on the way to site. Further updates will be provided as and when they become available.
Engineers are currently investigating an incident that is causing a loss of service on our VoIP platform. Further information will be provided as and when it becomes available.
We are currently receiving reports that a small number of DSL customers are experiencing difficulty when trying to send or receive emails. Our engineers are currently investigating and further information will be provided as and when it becomes available.
We are aware of an issue affecting a number of BT Ethernet circuits which connect to our Columbo core. Our initial diagnostics indicate the issue to be on BT’s hub and we have this raised with them already for investigation.
We will provide further information once it becomes available.
Thursday 6th July 2017 00:01 – 06:00
During the above window we will be replacing man.core.enta.net with a higher capacity router.
All services currently terminated in this location will be migrated on to the new hardware. This includes all customer transit, EWCS, wholesale hand-off and pseudowires terminating at this location.
All other traffic that would normally route through this location will take alternative paths.
We are currently investigating an issue affecting some of our Viper hosting platform.
Engineers are currently investigating and further updates will be provided when available.
Wednesday 05 July 2017 00:00 – 4:00
During the above window we will be moving some customers with EAD circuits directly connected to faraday.core from one switch to another. These customers will be contacted directly before the window and should only experience a small amount of downtime each. However as a precaution we would consider all customers terminating in Faraday at risk.
We are aware that our Bristol.core router has unexpectedly reloaded. Customers routed via this node will have temporarily lost service while the reload completed. Currently, we do not believe that the reload is related to the ongoing planned work. We will investigate the cause and will report as appropriate. Apologies for the inconvenience caused.
Wednesday 21 Jume 2017 21:00 – 23:59
During the above window we will be making configuration changes to the DSL platform to enable additional interconnects.
This work is not expected to be customer affecting.